# Integrated Sensing and Communications in Downlink FDD MIMO without CSI Feedback

### v1

@@ VI Simulation Results
(b) Multi-target case, targets are at -60°, 0°, 72°.
Figure 2: Beam pattern example for $N=8,K=4,M=4$.
@@
  • GPI-ISAC (no RS): As a recently proposed ISAC framework under imperfect CSIT [2], it considers the problem in (35); yet with conventional SDMA.
@@
VI-A Sensing Beam Shaping for the Desired Target
In Fig. 2, we plot the beam patterns to investigate the sensing performance for each technique (with $\nu=\nu_{\textrm{max}}$ for the GPI cases). This analysis includes both single-target and multi-target scenarios, with the desired targets specified in Fig. 2. First, the Radar-only technique introduced in [2] demonstrates the lowest sidelobe levels in both cases, resulting in the highest beam gain at the target angle. This is expected as these designs exclusively prioritize the sensing metric without considering the communication metric. In contrast, other ISAC-supported methods exhibit relatively higher sidelobe levels. Notably, our proposed method (GPI-ISAC-RS) achieves the lowest sidelobe levels while maintaining high beam gain for the target direction, suggesting that the method achieves the highest possible sensing performance compared to other approaches. Another notable observation is that when the error covariance matrix estimation is not applied in the proposed method, there is an increase in sidelobe levels and a degradation of beam gain at the target angle, leading to reduced sensing performance. This shows that, even though the error covariance matrix estimation is used mainly for the communication metric, it also impacts the sensing performance in ISAC systems where hardware resources are shared. Furthermore, the RS case exhibits improved maximum sensing performance compared to the RS case without RS. This can be attributed to the common message stream in the RS configuration, which, as seen in (51), contributes to improving the sensing performance for the sensing-centric regime.
@@ VI-C Performance per SNR
In Fig. 4-(a), we represent the ergodic sum SE per SNR. Firstly, we recall that the SDR-based algorithm without rank-1 approximation achieves the same SE as GPI-ISAC (no RS) while providing a lower MSE. However, this approach is not practical, as feasible beamforming vectors require the rank-1 approximation. It also reveals that GPI-ISAC (no RS) provides higher SE than the SDR-based algorithm after rank-1 approximation (SDR-Rank1) and RZF for $T_{\text{mse}}=-11$ dB, while closely meeting the MSE constraint, as shown in Fig. 4-(b). In this regard, we observe that GPI-ISAC-noRS significantly outperforms SDR-Rank1 in terms of SE while satisfying the MSE constraints. This gap is mainly due to the fact that SDR-Rank1 is a radar-oriented ISAC beamforming algorithm as identified in Fig. 4-(b). Moreover, it is observed that the proposed RS-based ISAC precoding method achieves significant SE gains while maintaining the same MSE as the GPI-ISAC-noRS approach. This suggests that the proposed RS method effectively mitigates the increased multi-user interference caused by CSIT inaccuracies in zero-feedback scenarios.
@@ VII Conclusion
This paper presents a precoding method for FDD MIMO ISAC systems that removes the need for channel feedback, thereby simultaneously supporting sensing and communication functionalities with low-latency characteristics. Specifically, we formulate a problem to maximize the sum SE under a beam pattern constraint based on MSE. By reconstructing the DL channel solely from UL training data, we mitigate feedback overhead and address imperfections in CSIT through the estimation of the error covariance matrix. In addition, to mitigate interference between sensing and communication functionality, we consider RSMA that is known to provide robustness toward both objectives. Subsequently, we optimize the RSMA precoder using the KKT conditions to jointly update the precoding vector and Lagrange multipliers, wherein the principle of NEPv is used to resolve the non-convexity. Our extensive simulation results demonstrate that our method offers precise and versatile beam pattern control and substantial SE gains, outperforming current techniques in the sensing-communication performances.
@@
$\triangleq\log_{2}(\zeta(\bar{\mathbf{p}},\nu))$ (85)

### v3

@@ VI Simulation Results
(b) Multi-target case, targets are at -60°, 0°, 72°
Sidelobe suppression (in dB scale)
Method Single-target Multi-target
GPI-Radar-Only -10.8 -10.56
GPI-ISAC-RS -10.2 -10.14
GPI-ISAC-RS (w/o $\hat{\mathbf{\Sigma}}_{k}$) -9.92 -3.06
GPI-ISAC-noRS -7.61 -2.58
SDR-Opt -9.73 -4.08
Figure 2: Normalized beam patterns and corresponding sidelobe suppression evaluations for $N=8,K=4,M=4$.
@@
  • GPI-ISAC (no RS): As a recently proposed ISAC framework under imperfect CSI [2], it considers the problem in (35); yet with conventional SDMA.
@@
To evaluate the effectiveness of the proposed approach under realistic conditions, we conduct simulations assuming the UL and DL carrier frequencies are set to $f_{\mathrm{c}}^{\mathrm{ul}}=7.25$ GHz and $f_{\mathrm{c}}^{\mathrm{dl}}=7.75$ GHz, respectively.
VI-A Sensing meam shaping for desired target
Fig. 2 illustrates the beam patterns used to evaluate sensing performance for each technique (with $\nu=\nu_{\textrm{max}}$ in GPI-based methods). We consider both single- and multi-target scenarios, with target directions indicated in the figure. The Radar-only approach introduced in [2] shows the lowest sidelobe levels and the highest beam gain at the target angle in both cases, as it fully prioritizes sensing without accounting for communication. In contrast, ISAC-based methods generally exhibit higher sidelobe levels. Among them, the proposed GPI-ISAC-RS achieves the lowest sidelobe levels while maintaining strong main-lobe gain, indicating superior sensing performance. When the ECM is not applied in our method, both sidelobe levels and beam gain degrade, leading to reduced sensing performance. This demonstrates that ECM estimation, though originally intended to improve communication metrics, also benefits sensing performance in ISAC systems by enabling more efficient resource sharing. Furthermore, the use of RS contributes to additional gains, as the common stream in (51) reinforces the sensing objective in the sensing-centric regime.
To complement the beam pattern analysis with quantitative metrics, the table below the figures presents the sidelobe suppression values for each method shown in Fig. 2. These values represent the dB difference between the main-lobe peak and the highest sidelobe, based on the original beam patterns already normalized in the dB scale. As the table shows, our ECM-aided GPI-ISAC-RS achieves the greatest suppression across both scenarios, aligning well with the visual trends in Fig. 2.
@@ VI-C Performance per SNR
In Fig. 4-(a), we present the ergodic sum spectral efficiency (SE) versus SNR for the proposed RSMA-based ISAC precoding and baseline methods. The SDR-based algorithm without rank-1 approximation achieves identical SE to GPI-ISAC (no RS) but with lower MSE, as shown in Fig. 4-(b). However, this approach is impractical, as feasible beamforming vectors require rank-1 approximation. Consequently, GPI-ISAC (no RS) outperforms the SDR-based algorithm with rank-1 approximation (SDR-Rank1) and regularized zero-forcing (RZF) in SE for $T_{\text{mse}}=-11$ dB, while closely satisfying the MSE constraint, as depicted in Fig. 4-(b). This performance gap arises because SDR-Rank1 prioritizes radar-oriented beamforming, limiting communication efficiency. In contrast, the proposed RSMA-based ISAC precoding achieves significant SE gains over GPI-ISAC (no RS) while maintaining equivalent MSE. RSMA’s advantage stems from its common stream, which mitigates multi-user and sensing-to-communication interference in zero-feedback scenarios with imperfect CSIT. By encoding common message parts into a shared stream, decodable by all users via SIC, RSMA aligns with radar beampattern requirements, reducing interference from sensing signals, as explained by [4]. This robustness, detailed in [28, 25], ensures superior SE under CSIT inaccuracies as evidenced in Fig. 4-(a).
@@ VI-D Detection probability
VI-E Convergence
In Fig. 6, we illustrate the numerical convergence behavior of the proposed framework. As shown in the figure, the algorithm generally converges within approximately $15-20$ iterations, while higher SNR levels require more iterations. We note that, as well observed in [25], higher SNR levels demand more delicate adjustments to the precoder due to the increased amount of interference, which delays convergence.
Remark 6.
(Computational complexity analysis) The proposed GPI-ISAC algorithm optimizes precoding with a complexity driven by the inversion of the block diagonal matrix ${\bf{\Omega}}(\bar{\bf{p}})$, comprising $(K+M)$ submatrices of size $N\times N$, requiring $\mathcal{O}(N^{3}(K+M))$ per iteration. Including the feasibility check, the total complexity is $\mathcal{O}(\xi_{\rm out}(LN(K+M)+\xi_{\rm in}N^{3}(K+M)))$, where $\xi_{\rm in}$ and $\xi_{\rm out}$ are inner and outer iterations, and $L$ is the number of constraints. For typical cases where $L\leq\xi_{\rm in}N^{2}$, this simplifies to $\mathcal{O}(\xi_{\rm out}\xi_{\rm in}N^{3}(K+M))$. In contrast, the QSDP-based approach in [3, 9] incurs $\mathcal{O}(K^{6.5}M^{6.5}\log(1/\epsilon))$, which scales poorly for large $M$. For massive MIMO-ISAC systems with numerous antennas, our method’s linear scaling in $K+M$ offers significant computational savings, crucial for low-latency applications.
Figure 6: Convergence behavior of the proposed GPI-ISAC-RS algorithm over iterations, showing ergodic sum SE versus iteration count for various channel realizations.
VI-F Different sensing metrics: SCNR
Even that the proposed GPI-ISAC-RS mainly considers the beam-pattern MSE as a sensing performance metric, it is also possible to extend it by incorporating different sensing metrics. We demonstrate that our method is also applicable when SCNR is used as a sensing metric. In sensing, a direct line-of-sight is often required, and signals reflected from non-target objects, termed clutter, degrade sensing accuracy [2]. The received radar echo signal is modeled as: ${\bf y}_{\rm R}={\bf G}_{\rm tar}{\bf x}+{\bf G}_{\rm cl}{\bf x}+{\bf n}_{\rm R}$, where ${\bf G}_{\rm tar}=\sum_{i=1}^{S}\beta_{i}^{\rm tar}{\bf a}_{\rm r}(\theta^{\rm tar}_{i}){\bf a}_{\rm t}^{\sf H}(\theta^{\rm tar}_{i})$ represents target reflections, ${\bf G}_{\rm cl}=\sum_{j=1}^{C}\beta_{j}^{\rm cl}{\bf a}_{\rm r}(\theta^{\rm cl}_{j}){\bf a}_{\rm t}^{\sf H}(\theta^{\rm cl}_{j})$ denotes clutter, and ${\bf n}_{\rm R}\sim\mathcal{CN}({\bf 0},\sigma^{2}_{\rm R}{\bf I}_{N})$ is noise. Here, $S$ and $C$ are the number of targets and clutters, $\beta_{i}^{\rm tar}$, $\beta_{j}^{\rm cl}$ are reflection coefficients, and $\theta_{i}^{\rm tar}$, $\theta_{j}^{\rm cl}$ are directions, with corresponding array response vectors ${\bf a}_{\rm t}(\theta)$, ${\bf a}_{\rm r}(\theta)$. Based on this, the SCNR is derived as
$\gamma_{\rm R}(\bar{\bf{p}})$ $=\frac{\bar{\bf p}^{\sf H}({\bf I}_{K+M}\otimes{\bf G}_{\rm tar})^{\sf H}({\bf I}_{K+M}\otimes{\bf G}_{\rm tar})\bar{\bf p}}{\bar{\bf p}^{\sf H}({\bf I}_{K+M}\otimes{\bf G}_{\rm cl})^{\sf H}({\bf I}_{K+M}\otimes{\bf G}_{\rm cl})\bar{\bf p}+\frac{N}{P}\sigma^{2}_{\rm R}}$ (84)
$=\frac{\bar{\bf p}^{\sf H}\bar{\bf{G}}_{\rm tar}\bar{\bf p}}{\bar{\bf p}^{\sf H}\bar{\bf{G}}_{\rm cl}\bar{\bf p}},$ (85)
which effectively incorporates clutter impact, making it suitable for radar tracking with known target and clutter positions. Based on this, the SCNR-constrained ISAC optimization problem is formulated based on $\mathscr{P}_{2}$ as
$\mathscr{P}_{\text{scnr}}:$ $\mathop{\textrm{missing}}{maximize}_{\bar{\mathbf{p}}}\ \bar{R}^{\textrm{LSE}}_{\mathrm{c}}(\bar{\mathbf{p}})+\sum_{k=1}^{K}\bar{R}_{k}(\bar{\mathbf{p}})$ (86)
subject to $\textrm{subject to }\frac{\bar{\bf p}^{\sf H}\bar{\bf G}_{\rm tar}\bar{\bf p}}{\bar{\bf p}^{\sf H}\bar{\bf G}_{\rm cl}\bar{\bf p}}\geq T_{\rm scnr}.$ (87)
To solve this, we adopt the similar approach to the proposed GPI-ISAC-RS. We obtain the first-order KKT condition of the problem $\mathscr{P}_{\text{scnr}}$ as ${\bf\Upsilon}(\bar{\bf p})\bar{\bf p}={\bf\Xi}(\bar{\bf p})\eta(\bar{\bf p})\bar{\bf p}$. Here, ${\bf\Upsilon}(\bar{\bf p})$ and ${\bf\Xi}(\bar{\bf p})$ are obtained by replacing the sensing-related term in $\mathbf{L}(\bar{\mathbf{p}},\mu)$ (1) and $\mathbf{R}(\bar{\mathbf{p}},\mu)$ (68) by
$\mu\frac{\bar{\bf G}_{\rm tar}}{\bar{\bf p}^{\sf H}\bar{\bf G}_{\rm tar}\bar{\bf p}},\;\mu\frac{\bar{\bf G}_{\rm cl}}{\bar{\bf p}^{\sf H}\bar{\bf G}_{\rm cl}\bar{\bf p}},$ (88)
respectively, where $\mu$ indicates the Lagrangian multiplier. Building on this, we find the principal eigenvector that solves the KKT condition by slightly modifying the GPI-ISAC-RS. Due to space limitations, we omit the detailed development.
In Fig. 7-(a), we observe that the proposed GPI-ISAC-RS consistently outperforms the method in [2] in all SCNR targets. Clearly, these gains come from the use of RSMA combined with the proposed ECM estimation, which effectively suppresses the interference caused by DL CSI reconstruction without no direct feedback. It is notable that the sensing beam pattern is suitably formed by incorporating the target and clutter directions as shown in Fig. 7-(b).
(a) Ergodic sum SE per target SCNR
(b) Sensing beam pattern with target SCNR = 15 dB
Figure 7: Ergodic sum SE and sensing beam pattern when considering SCNR as a new sensing metric. $N=8,K=4,M=4,$ and $\eta_{k,\ell}=0.9\ \forall(k,\ell)$.
VI-G Large scale antenna regime
We compare the performance of the proposed method with that of previous ISAC precoding method [2] as increasing the number of transmit antennas. We fix the number of users to be half the number of antennas (i.e., $K=N/2$), and plot the average SE per user (instead of the sum SE). As shown in Fig. 8-(a), the proposed method consistently outperforms GPI-ISAC across all antenna configurations. The relative gain becomes less pronounced as the number of antennas increases. This is because as $N$ grows, the channel vectors become more orthogonal [24], thereby reducing the interference. This diminishes the relative benefit of the proposed precoding. In Fig. 8-(b), as $N$ increases, the sensing beam becomes sharper at the target direction while the gain at non-target directions is well suppressed. This confirms that the proposed method effectively exploits the increased spatial DoF in large scale antenna regime to enhance radar resolution.
(a) Ergodic sum SE per number of antennas
(b) Sensing beam pattern with target MSE = -9 dB. The targets are located at -50°and 50°.
Figure 8: Ergodic sum SE and sensing beam pattern in large scale antenna regimes. $K=N/2$ is assumed.
@@ VII Conclusion
This paper presents a precoding method for FDD MIMO ISAC systems that removes the need for channel feedback, thereby simultaneously supporting sensing and communication functionalities with low-latency characteristics. In particular, we aim to maximize the ergodic sum spectral efficiency while satisfying a beam pattern constraint based on the MSE. By reconstructing the DL channel solely from UL training data, we reduce feedback overhead and compensate for CSI imperfections by estimating the ECM. In addition, to mitigate interference between sensing and communication functionality, we consider RSMA that is known to provide robustness toward both objectives. Subsequently, we optimize the RSMA precoder using the KKT conditions to jointly update the precoding vector and Lagrange multipliers, wherein the principle of NEPv is used to resolve the non-convexity. Our extensive simulation results demonstrate that our method offers precise and versatile beam pattern control and substantial SE gains, outperforming current techniques in the sensing-communication performances.
@@
$\triangleq\log_{2}(\zeta(\bar{\mathbf{p}},\nu)).$ (90)